Grade 12th passMechanics

An object thrown vertically up from the ground passes the height 5m twice in an interval of 10 sec. What is its time of flight? Will it be right to take 5 m as the highest point in this question? Please explain.

Aabha Misra
8 Years agoGrade 12th pass

1 Answer

Sandeep Bhamoo
8 Years ago
Take this height of 5m as it`s lowest point in this interval of 10s ,hence it reaches max. height at 5s and at max. height it`s velocity is zero since v=u`+at. 0=u`-10•5. u`=50m/s , it is the velocity of object at a height of 5m and since v² -u²= 2as. v=u`,u=u°-initial vel. at ground s=5m. putting these values (u°)²=3500. u°=10√35m/s. and time of flight = 2u°/a=(20√35)/(20)= √35s. No it is not right to take 5m as heighest point.